Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probabilty question: where m i wrong??

Expert replies
by him1985 » Sat Jan 28, 2012 10:23 pm
Q. A certain deck of cards contains 2 blue cards, 2 red cards, 2 yellow cards, and 2 green cards. If two cards are randomly drawn from the deck, what is the probability that they will both are not blue?

A. 15/28

B. 1/4

C. 9/16

D. 1/32

E. 1/16


We can have two approches here:
(1). direct find probabilty that both balls are not blue.
6/8 * 5/7 = 15/28

(2). We can find the prob. that both balls are blue and can substract from 1

1- (2/8*1/7) = 27/28

I am getting two different answer but according to probability rule i should get answer same in both way. Please help....
Himanshu Chauhan
Join the discussion
Source: — Problem Solving |

by rijul007 » Sat Jan 28, 2012 11:00 pm
him1985 wrote:Q. A certain deck of cards contains 2 blue cards, 2 red cards, 2 yellow cards, and 2 green cards. If two cards are randomly drawn from the deck, what is the probability that they will both are not blue?

A. 15/28

B. 1/4

C. 9/16

D. 1/32

E. 1/16


We can have two approches here:
(1). direct find probabilty that both balls are not blue.
6/8 * 5/7 = 15/28

(2). We can find the prob. that both balls are blue and can substract from 1

1- (2/8*1/7) = 27/28

I am getting two different answer but according to probability rule i should get answer same in both way. Please help....
Analyze the appraoches yourself, and answer this ques for each appraoch
does the computed probability include the selection of one blue and one green?

Did you see the difference?
Join the discussion

by LalaB » Sun Jan 29, 2012 1:28 am
him1985,

the probability that both are not blue=1-(probability that one or both are blue)

the probability that they will both are not blue=1-(2C2+2C1*6C1)/8C2=1-13/28=15/28

One more approach-

6C2/8C2=15/28
Join the discussion

by ronnie1985 » Sun Jan 29, 2012 10:59 am
No blue = 6C2
2 balls = 8C2
P = 6c2/8C2 = 15/28.
Follow your passion, Success as perceived by others shall follow you
Join the discussion

by him1985 » Sun Jan 29, 2012 6:13 pm
LalaB wrote:him1985,

the probability that they will both are not blue=1-(2C2+2C1*6C1)/8C2=1-13/28=15/28
LalaB: Can you please elaborate this calculation. Probability that both are blue =(2C2+2C1*6C1)/8C2
Himanshu Chauhan
Join the discussion

by [email protected] » Sun Jan 29, 2012 9:06 pm
him1985 wrote:
LalaB wrote:him1985,

the probability that they will both are not blue=1-(2C2+2C1*6C1)/8C2=1-13/28=15/28
LalaB: Can you please elaborate this calculation. Probability that both are blue =(2C2+2C1*6C1)/8C2
I can elaborate.
selecting 2 blue out of 2 blue =2C2=1
selecting 1 blue out of 1 blue and selecting any other out of 6 different=2C1*6C1=12

So Prob is =1-13/8C2=1-13/28=15/28
Join the discussion